OpenStudy (anonymous):

Why is the product of the gradients of 2 straight line perpendicular to each other is always equal to -1 ?? Please prove it......?

OpenStudy (anonymous):

m1=1/m2 you can use two rules (straight line) to prove it. 

OpenStudy (anonymous):

Give me a hint?

OpenStudy (anonymous):

you can write equation of the lines perpedicular to each other in this form: ax+by=0 -bx+ay=0 this is a hint....:)

OpenStudy (anonymous):

Can you prove it through trigonometry?

OpenStudy (anonymous):

@myko @shirleyx Can you prove it through trigonometry?

OpenStudy (experimentx):

slope = \( \tan(\theta) \) another slope = \( \tan(\theta - 90) = -\cot \theta \)

OpenStudy (anonymous):

\[\cos \alpha x+\sin \alpha y =0\] \[-\sin \alpha x + \cos \alpha y =0\]

OpenStudy (anonymous):

grad f1 = (cosa -sinax,sina+cosay) grad f2 = (-sina-cosax,cosa-sinay)